# Stong order $1$ adaptive approximation of jump-diffusion SDEs with discontinuous drift

### v1

@@ 4 Convergence of the transformation-based doubly-adaptive quasi-Milstein scheme
${\mathbb{E}}\Big{[}\sup_{t\in[0,T]}|X_{t}-G^{-1}(Z^{(\delta)}_{t})|^{p}\Big{]}^{\frac{1}{p}}\leq c_{12}\delta.$ (4.7)
@@ Appendix A Additional proofs
$\tau_{n+1}=$ $\big{(}\tau_{n}+h^{\delta}(Z^{(\delta)}_{\tau_{n}})\big{)}\wedge\big{(}\tau_{n}^{\delta}+\inf\{s\geq 0:N_{\tau_{n}+s}-N_{\tau_{n}}\geq 0\}\big{)}\wedge T,$ (A.1)
which implies that $\tau_{n+1}$ can be expressed as a measurable function $f$ of $\tau_{n}$, $Z^{(\delta)}_{\tau_{n}}$ and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$. Using this we obtain
@@
$\leq{\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big{]}.$
Applying [15, p. 33] we get
@@
$\leq{\mathbb{E}}\Big{[}{\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big{|}\mathbb{F}_{\tau_{n}}\Big{]}\Big{]}$
$={\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}{\mathbb{E}}\Big{[}\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\Big{|}\mathbb{F}_{\tau_{n}}\Big{]}\Big{]}$
@@
$\quad\quad\quad\cdot{\mathbb{E}}\Big{[}\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(z,\widetilde{z},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\Big{|}\mathbb{F}_{\tau_{n}}\Big{]}\Big{|}_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big{]}.$
By Lemma 3.3, $(W_{\tau_{n}+s}-W_{\tau_{n}})_{s\geq 0}$ and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$ are independent of $\mathbb{F}_{\tau_{n}}$ as well as $(W_{\tau_{n}+s}-W_{\tau_{n}})_{s\geq 0}$ is independent of $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$. This, Lemma 3.17, and [15, p. 33] imply
${\mathbb{E}}\big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}|W_{t}-W_{\tau_{n}}|^{q}\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{u}=k\}}\big{]}$ (A.4)
$\leq c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}{\mathbb{E}}\Big{[}\mathds{1}_{\{f(z,\widetilde{z}(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\Big{]}\Big{|}_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big{]}$
$=c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big{]}$
$=c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{u}=k\}}\Big{]}.$
Further,
@@
$=\sum_{k=0}^{\infty}\sum_{n=0}^{\infty}{\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}|W_{t}-W_{\tau_{n}}|^{q}\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{u}=k\}}\Big{]}$
$\leq\sum_{k=0}^{\infty}\sum_{n=0}^{\infty}c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big{[}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}}\big{|}^{p}\big{)}\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{u}=k\}}\Big{]}$
@@
By Hölder’s inequality it is enough to prove the claim for $p\in{\mathbb{N}}\setminus\{1\}$. For the first summand it is enough to consider instead
@@
By (3.9) and the fact that $N_{\tau_{n}}-N_{\tau_{n-1}}\in\{0,1\}$ we get that for all $n$, $k\in{\mathbb{N}}$ there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that
@@
Further by Lemma 3.3 and Lemma 3.17 we obtain for all $q\in{\mathbb{N}}$,
@@
By (3.8) and (LABEL:FM6) there exists constants $\widetilde{c}_{2},\widetilde{c}_{3},\widetilde{c}_{4}\in(0,\infty)$ such that for all $k,n\in{\mathbb{N}}$,
@@
Together with (LABEL:FM4) we obtain that there exist constants $\widetilde{c}_{5},\widetilde{c}_{6}\in(1,\infty)$ such that
@@
Using (A.10) recursively we get that there exist constants $\widetilde{c}_{7},\widetilde{c}_{8}\in(0,\infty)$ such that
@@
Analogously we obtain using recusively (A.11), and (LABEL:FM6) that there exists a constant $\widetilde{c}_{9}\in(0,\infty)$ such that
@@
Combining (LABEL:FM3) and (LABEL:FM9) and using that $\widetilde{c}_{5}\in(1,\infty)$ we obtain
@@
$\leq\frac{T\,\widetilde{c}_{8}\,\widetilde{c}_{5}^{\,M+2}}{\widetilde{c}_{5}-1}\sum_{k=0}^{\infty}{\mathbb{P}}(N_{T}=k)\,\widetilde{c}_{5}^{\,k}\leq\frac{T\,\widetilde{c}_{8}\,\widetilde{c}_{5}^{\,M+2}}{\widetilde{c}_{5}-1}\exp(\lambda T(\widetilde{c}_{5}-1)).$
For the second summand we use that $N_{T}$ has finite $p$-th moments and calculate similar to (LABEL:FM10) using the fact that $\widetilde{c}_{6}\in(1,\infty)$,
@@
$\leq\frac{\widetilde{c}_{10}\widetilde{c}_{9}\widetilde{c}_{6}^{\,M+1}}{\widetilde{c}_{6}-1}\exp(\lambda T(\widetilde{c}_{6}-1)),$
where $\widetilde{c}_{10}$ is the $(p-1)$-th moment of a Poisson distributed random variable with parameter $\widetilde{c}_{6}\lambda T$. ∎
@@
By Hölder’s inequality it is enough to prove the claim for $p\in{\mathbb{N}}\setminus\{1\}$. Recalling equation (3.11) we obtain for all $s\in[0,T]$,
@@
Now we estimate each summand separately. For the first one we apply Jensen’s inequality, and Lemma 3.8 to obtain
@@
For the second summand we apply Lemma 3.6 to show that there exists $\hat{c}\in(0,\infty)$ such that
@@
Using Lemma 3.7 and Lemma 3.8 we obtain that there exist $c_{\delta},c_{W_{p}}\in(0,\infty)$ such that
@@
For the last summand, we recall that $\nu_{i}$ denotes the $i$-th jump time of $N$ and use Lemma 3.8 to obtain that $\delta\in(0,\delta_{0}]$, and $M=\Big{\lceil}\frac{T}{\delta^{2}\log^{4}(\delta^{-1})}\Big{\rceil}$ there exists $c_{\delta}\in(0,\infty)$ such that
@@
$\leq 2^{p-1}c_{\widetilde{\rho}}^{p}\,{\mathbb{E}}\Big{[}N_{s}^{p-1}\sum_{i=1}^{N_{s}}\big{(}1+\big{|}Z^{(\delta)}_{\nu_{i}-}\big{|}^{p}\big{)}\Big{]}\leq 2^{p-1}c_{\widetilde{\rho}}^{p}\,{\mathbb{E}}\Big{[}N_{T}^{p-1}\sum_{n=0}^{N_{T}+M}\big{(}1+\big{|}Z^{(\delta)}_{\tau_{n}-}\big{|}^{p}\big{)}\Big{]}\leq 2^{p-1}c_{\widetilde{\rho}}^{p}c_{\delta}.$
Plugging (LABEL:ME2), (LABEL:ME16), and (LABEL:ME17) in (LABEL:ME1) leads that for all $s\in[0,T]$, all $p\in{\mathbb{N}}$, and all $\delta\in(0,\delta_{0}]$, $M=\Big{\lceil}\frac{T}{\delta^{2}\log^{4}(\delta^{-1})}\Big{\rceil}$, ,
@@
Next we change our previous estimates with the goal to apply Gronwall’s inequality. We again start with (LABEL:ME1). We change the estimate (LABEL:ME2) to
@@
We change the estimate in (LABEL:ME16) to
@@
For the last summand of (LABEL:ME1) by Lemma 3.6 there exists $\hat{c}\in(0,\infty)$ such that
@@
Combining (LABEL:ME1), (LABEL:ME19), (LABEL:ME24), and (LABEL:ME25) we get that there exists a constants $\widetilde{c}_{1}\in(0,\infty)$ such that
@@
As $[0,T]\ni s\mapsto{\mathbb{E}}\big{[}\sup_{v\in[0,s]}\big{|}Z^{(\delta)}_{v}\big{|}^{p}\big{]}$ is s finite valued, Borel measurable function we obtain using Gronwall’s inequality that there exists a constant $c_{1}\in(0,\infty)$ such that
@@
Next we observe that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that
@@
Applying Lemma 3.17 to the second and third summand and using (LABEL:ME27) we get that there exists a constant $\widetilde{c}_{3}\in(0,\infty)$ such that
@@
For the last statement assume $t\in[0,T-\Delta]$. By Lemma 3.6 there exists a constant $\hat{c}\in(0,\infty)$ such that
@@
For the first summand we estimate
@@
For the second summand we obtain by Lemma 3.7 that there exists $c_{W_{p}}\in(0,\infty)$ such that
@@
For the third summand we get
@@
Plugging (A.30), (A.31), and (A.32) into (A.29) yields that there exists a constant $\widetilde{c}_{4}\in(0,\infty)$ such that
@@
Recall that $\tau_{n+1}$ can be expressed as a measurable function $f$ of $\tau_{n}$, $Z^{(\delta)}_{\tau_{n}}$ and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$. Hence for all $n\in{\mathbb{N}}$ and $k\in{\mathbb{N}}$,
@@
$\quad\quad\quad\quad\quad\quad\quad\cdot\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big{]}.$
By [15, p. 33],
@@
$\quad\quad\quad{\mathbb{E}}\Big{[}\sup_{s\in[0,\delta^{2}\log^{4}(\delta^{-1})]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\Big{|}\mathbb{F}_{\tau_{n}}\Big{]}\Big{]}$
@@
$\quad\quad\quad\cdot{\mathbb{E}}\Big{[}\sup_{s\in[0,\delta^{2}\log^{4}(\delta^{-1})]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(z,\widetilde{z},(N_{s+\tau_{n}}-N_{\tau_{n}}))_{s\geq 0})>t\}}\Big{|}\mathbb{F}_{\tau_{n}}\Big{]}\Big{|}_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big{]}.$
By Lemma 3.3, Lemma 3.17, and [15, p. 33] we get

### v2

@@ 4 Convergence of the transformation-based doubly-adaptive quasi-Milstein scheme
${\mathbb{E}}\Big[\sup_{t\in[0,T]}|X_{t}-G^{-1}(Z^{(\delta)}_{t})|^{p}\Big]^{\frac{1}{p}}\leq c_{12}\delta;$ (4.7)
@@ Appendix A Additional proofs
$\tau_{n+1}=$ $\big(\tau_{n}+h^{\delta}(Z^{(\delta)}_{\tau_{n}})\big)\wedge\big(\tau_{n}+\inf\{s\geq 0:N_{\tau_{n}+s}-N_{\tau_{n}}\geq 0\}\big)\wedge T,$ (A.1)
which implies that $\tau_{n+1}$ can be expressed as a measurable function $f$ of $\tau_{n}$, $Z^{(\delta)}_{\tau_{n}}$ and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$. Using this we obtain for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
$\leq{\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big].$
Applying [15, p. 33] we get for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
$\leq{\mathbb{E}}\Big[{\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big|\mathbb{F}_{\tau_{n}}\Big]\Big]$
$={\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}{\mathbb{E}}\Big[\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\Big|\mathbb{F}_{\tau_{n}}\Big]\Big]$
@@
$\quad\quad\quad\cdot{\mathbb{E}}\Big[\sup_{s\in[0,\delta]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(z,\widetilde{z},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\Big|\mathbb{F}_{\tau_{n}}\Big]\Big|_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big$].
By Lemma 3.3, $(W_{\tau_{n}+s}-W_{\tau_{n}})_{s\geq 0}$ and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$ are independent of $\mathbb{F}_{\tau_{n}}$ as well as $(W_{\tau_{n}+s}-W_{\tau_{n}})_{s\geq 0}$ is independent of $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$. This, Lemma 3.17, and [15, p. 33] imply for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
${\mathbb{E}}\big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)|W_{t}-W_{\tau_{n}}|^{q}\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{\tau_{n}}=k\}}\big]$ (A.4)
$\leq c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}{\mathbb{E}}\Big[\mathds{1}_{\{f(z,\widetilde{z},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\Big]\Big|_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big$]
$=c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big]$
$=c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big].$
Further, for all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
$=\sum_{k=0}^{\infty}\sum_{n=0}^{\infty}{\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)|W_{t}-W_{\tau_{n}}|^{q}\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big]$
$\leq\sum_{k=0}^{\infty}\sum_{n=0}^{\infty}c_{W_{q}}\delta^{\frac{q}{2}}{\mathbb{E}}\Big[\big(1+\big|Z^{(\delta)}_{\tau_{n}}\big|^{p}\big)\mathds{1}_{[\tau_{n},\tau_{n+1})}(t)\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big]$
@@
By Hölder’s inequality it is enough to prove the claim for $p\in[2,\infty)\cap{\mathbb{N}}$. For the first summand it is enough to consider instead for all $\delta\in(0,\delta_{0}]$,
@@
By (3.9) and the fact that $N_{\tau_{n}}-N_{\tau_{n-1}}\in\{0,1\}$ we get that for all $n$, $k\in{\mathbb{N}}$ there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,
@@
Further by Lemma 3.3 and Lemma 3.17 we obtain for all $q\in{\mathbb{N}}$, all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
By (3.8) and (LABEL:FM6) there exists constants $\widetilde{c}_{2},\widetilde{c}_{3},\widetilde{c}_{4}\in(0,\infty)$ such that for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
Together with (LABEL:FM4) we obtain that there exist constants $\widetilde{c}_{5},\widetilde{c}_{6}\in(1,\infty)$ such that for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
Using (A.10) recursively we get that there exist constants $\widetilde{c}_{7},\widetilde{c}_{8}\in(0,\infty)$ such that for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
Analogously we obtain using recusively (A.11), and (LABEL:FM6) that there exists a constant $\widetilde{c}_{9}\in(0,\infty)$ such that for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
Combining (LABEL:FM3) and (LABEL:FM9) and using that $\widetilde{c}_{5}\in(1,\infty)$ we obtain for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
$\leq\frac{T\,\widetilde{c}_{8}\,\widetilde{c}_{5}^{\,M+2}}{\widetilde{c}_{5}-1}\sum_{k=0}^{\infty}{\mathbb{P}}(N_{T}=k)\,\widetilde{c}_{5}^{\,k}=\frac{T\,\widetilde{c}_{8}\,\widetilde{c}_{5}^{\,M+2}}{\widetilde{c}_{5}-1}\exp(\lambda T(\widetilde{c}_{5}-1)).$
For the second summand we use that $N_{T}$ has finite $p$-th moments and calculate similar to (LABEL:FM10) using the fact that $\widetilde{c}_{6}\in(1,\infty)$, we obtain for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$ and all $\delta\in(0,\delta_{0}]$,
@@
$=\frac{\widetilde{c}_{9}\,\widetilde{c}_{6}^{\,M+1}}{\widetilde{c}_{6}-1}\sum_{k=0}^{\infty}k^{p-1}\frac{(\lambda T\widetilde{c}_{6})^{k}}{k!}e^{-\lambda T\widetilde{c}_{6}}e^{\lambda T(\widetilde{c}_{6}-1)}=\frac{\widetilde{c}_{10}\widetilde{c}_{9}\widetilde{c}_{6}^{\,M+1}}{\widetilde{c}_{6}-1}\exp(\lambda T(\widetilde{c}_{6}-1)),$
where $\widetilde{c}_{10}$ is the (p1)(p-1)-th moment of a Poisson distributed random variable with parameter $\widetilde{c}_{6}\lambda T$. ∎
@@
By Hölder’s inequality it is enough to prove the claim for $p\in[2,\infty)\cap{\mathbb{N}}$. Recalling equation (3.11) we obtain for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Now we estimate each summand separately. For the second one we apply Jensen’s inequality, and Lemma 3.8 to obtain for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
For the third summand we apply Lemma 3.6 to show that there exists $\hat{c}\in(0,\infty)$ such that for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Using Lemma 3.7 and Lemma 3.8 we obtain that there exist $c_{\delta},c_{W_{p}}\in(0,\infty)$ such that for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
For the last summand, we recall that $\nu_{i}$ denotes the $i$-th jump time of $N$ and use Lemma 3.8 to obtain that for all $\delta\in(0,\delta_{0}]$ there exists $c_{\delta}\in(0,\infty)$ such that for all $s\in[0,T]$,
@@
$\leq 2^{p-1}c_{\widetilde{\rho}}^{p}\,{\mathbb{E}}\Big[N_{s}^{p-1}\sum_{i=1}^{N_{s}}\big(1+\big|Z^{(\delta)}_{\nu_{i}-}\big|^{p}\big)\Big]\leq 2^{p-1}c_{\widetilde{\rho}}^{p}\,{\mathbb{E}}\Big[N_{T}^{p-1}\sum_{n=0}^{N_{T}+M}\big(1+\big|Z^{(\delta)}_{\tau_{n}-}\big|^{p}\big)\Big]\leq 2^{p-1}c_{\widetilde{\rho}}^{p}c_{\delta},$
where $M=\Big\lceil\frac{T}{\delta^{2}\log^{4}(\delta^{-1})}\Big\rceil$. Plugging (LABEL:ME2), (LABEL:ME16), and (LABEL:ME17) in (LABEL:ME1) leads that for all $s\in[0,T]$, all $p\in{\mathbb{N}}$, and all $\delta\in(0,\delta_{0}]$,
@@
Next we change our previous estimates with the goal to apply Gronwall’s inequality. We again start with (LABEL:ME1). We change the estimate (LABEL:ME2) for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$ to
@@
We change the estimate in (LABEL:ME16) for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$ to
@@
For the last summand of (LABEL:ME1) by Lemma 3.6 there exists $\hat{c}\in(0,\infty)$ such that for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Combining (LABEL:ME1), (LABEL:ME19), (LABEL:ME24), and (LABEL:ME25) we get that there exists a constants $\widetilde{c}_{1}\in(0,\infty)$ such that for all $s\in[0,T]$ and all $\delta\in(0,\delta_{0}]$
@@
As $0,T]\ni s\mapsto{\mathbb{E}}\big[\sup_{v\in[0,s]}\big|Z^{(\delta)}_{v}\big|^{p}\big]$ is a finite valued, Borel measurable function we obtain using Gronwall’s inequality that there exists a constant $c_{1}\in(0,\infty)$ such that for all $t\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Next we observe that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that for all $t\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Applying Lemma 3.17 to the second and third summand and using (LABEL:ME27) we get that there exists a constant $\widetilde{c}_{3}\in(0,\infty)$ such that for all $t\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
For the last statement by Lemma 3.6 there exists a constant $\hat{c}\in(0,\infty)$ such that for all $t\in[0,T-\Delta]$ and all $\delta\in(0,\delta_{0}]$,
@@
For the first summand we estimate for all $u\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
For the second summand we obtain by Lemma 3.7 that there exists $c_{W_{p}}\in(0,\infty)$ such that for all $u\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
For the third summand we get for all $u\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Plugging (A.30), (A.31), and (A.32) into (A.29) yields that there exists a constant $\widetilde{c}_{4}\in(0,\infty)$ such that for all $t\in[0,T-\Delta]$ and all $\delta\in(0,\delta_{0}]$,
@@
Recall that $\tau_{n+1}$ can be expressed as a measurable function $f$ of $\tau_{n}$, $Z^{(\delta)}_{\tau_{n}}$ and $(N_{\tau_{n}+s}-N_{\tau_{n}})_{s\geq 0}$. Hence for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
$\quad\quad\quad\quad\quad\quad\quad\cdot\mathds{1}_{\{\tau_{n}\leq t\}}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\mathds{1}_{\{N_{\tau_{n}}=k\}}\Big].$
By [15, p. 33], for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
$\quad\quad\quad{\mathbb{E}}\Big[\sup_{s\in[0,\delta^{2}\log^{4}(\delta^{-1})]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(\tau_{n},Z^{(\delta)}_{\tau_{n}},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\Big|\mathbb{F}_{\tau_{n}}\Big]\Big]$
@@
$\quad\quad\quad\cdot{\mathbb{E}}\Big[\sup_{s\in[0,\delta^{2}\log^{4}(\delta^{-1})]}|W_{s+\tau_{n}}-W_{\tau_{n}}|^{q}\mathds{1}_{\{f(z,\widetilde{z},(N_{s+\tau_{n}}-N_{\tau_{n}})_{s\geq 0})>t\}}\Big|\mathbb{F}_{\tau_{n}}\Big]\Big|_{z=\tau_{n},\widetilde{z}=Z^{(\delta)}_{\tau_{n}}}\Big$].
By Lemma 3.3, Lemma 3.17, and [15, p. 33] we get for all $n\in{\mathbb{N}}$, all $k\in{\mathbb{N}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,